OpenStudy (anonymous):

what is the solution of the equation 6y-2(y+1)=3(y-2)+6

OpenStudy (anonymous):

6y - 2y + 2 = 3y - 6 + 6 4y + 2 = 3y 4y - 3y = -2

OpenStudy (kamibug):

1. Distribute 6y -2y +2 = 3y -6 +6 2. Simplify by adding Like Terms 4y +2 = 3y 3. Isolate the Variable y = -2